OpenStudy (lilsis76):

express sin(17pi/12) in terms of rational numbers and radicals. how can i do this :/

OpenStudy (lilsis76):

@anikay think u can help me with this one?

OpenStudy (anikay):

remember what I said about pi/6?

OpenStudy (lilsis76):

er.....hold on, u said....pi//6 is 30 deg.

OpenStudy (lilsis76):

@anikay

OpenStudy (anikay):

ok, so, what would 12 be? then?

OpenStudy (lilsis76):

let me draw it reallyfast

OpenStudy (anikay):

you're going to need to know your half angle formulas for this

OpenStudy (lilsis76):

ugh. i forgot it :/

OpenStudy (anikay):

well, sucks, because it's gonna be a 75 degree reference angle :P

OpenStudy (lilsis76):

hold on im looking up the formula

OpenStudy (lilsis76):

cos(2a)=cos^2 - sine^2 = 2cos^2 - 1 = 1 -2 sin^2

OpenStudy (lilsis76):

right? @anikay

OpenStudy (anikay):

awww D: no that's double angle

OpenStudy (lilsis76):

okay hold on let me check again

OpenStudy (anikay):

at least, the cos(2a) is

OpenStudy (lilsis76):

okay i got it, sin Q = 2 sin Q/2 cos Q/2, right?

OpenStudy (anikay):

uh, here

OpenStudy (lilsis76):

haha im not good at math seriously ugh, this is horrible. im sorry if im wasting your time

OpenStudy (anikay):

meh tryin to hit multiple problems

OpenStudy (lilsis76):

haha thank you

OpenStudy (anikay):

ok so its

OpenStudy (anikay):

\[\cos ^2 \frac{ \theta }{ 2 } = \frac{ 1+\cos(\theta) }{ 2 }\]

OpenStudy (lilsis76):

hm...okay, do i ....just plug in the 17pi/12?

OpenStudy (anikay):

uh....im tired xD

OpenStudy (anikay):

imma go with no

OpenStudy (anikay):

imma say plug in 150

OpenStudy (lilsis76):

okay :/ then ill skip it

OpenStudy (lilsis76):

why 150?

OpenStudy (anikay):

cuz 150/2 = 75 xD or you could try 17pi/6

OpenStudy (lilsis76):

haha where is the 17 pi coming froM?!

OpenStudy (anikay):

isn't that your original problem?

OpenStudy (anikay):

but multiplied by 2?

OpenStudy (lilsis76):

IT IS!!! haha i finally saw it haha

OpenStudy (lilsis76):

okay i have one more problem and its my last.

OpenStudy (anikay):

yesh?

OpenStudy (anikay):

I should sleep soon.... is this a better pastime than League of Legends? xD